### v2

# A projector–rank partition theorem for exact degrees of freedom in experimental design

Abstract
Degrees of freedom ($\mathrm{df}$) allocation becomes opaque once blocking, nesting, fractionation, or unequal replication disturb the balanced one-stratum classical ANOVA. We resolve this ambiguity with a $\mathrm{df}$ partition theorem. For $N$ observations and a set $\mathcal{S}$ that indexes all randomization strata in the experiment, let $\mathbf{P}_{s}$ denote the idempotent projector that averages observations within randomization stratum $s\in\mathcal{S}$ and $\mathcal{C}_{\overline{E}}$ the factorial-contrast subspace attached to the possibly aliased effect $\overline{E}$. We prove the rank identity $N-1=\sum_{s}\sum_{\overline{E}}\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$, which simultaneously diagonalizes all strata and all effects, yielding exact integer $\mathrm{df}$ for any fixed-random mixture, block structure, or replication pattern. It results in closed-form $\mathrm{df}$ tables for unbalanced designs previously reliant on numerical approximations. We introduce aliasing measures such as the $\mathrm{df}$-retention ratio $\rho(\overline{E})$, deficiency $\delta(\overline{E})$ and the variance-inflation index $\alpha(\overline{E})$ that extend Box–Hunter’s resolution concept to multi-stratum and incomplete designs, and supply immediate diagnostics of information loss. Classical results such as Cochran’s identity when $\lvert\mathcal{S}\rvert=1$ and balanced fractions, Yates’ $\mathrm{df}$ for balanced split-plots, and resolution-$R$ when $\rho(\overline{E})=0$ emerge as corollaries. The method yields calibrated error rates, boosts power by up to 60%, and outpaces bootstrap approximations.
Keywords: Stratified ANOVA, projection matrices, alias structure analysis.
1 Introduction
Modern experimentation, whether in agriculture, manufacturing, or the life sciences, rarely conforms to the tidy assumptions that powered the early triumphs of analysis of variance (ANOVA) pioneered by Fisher, (1935). Blocking against environmental heterogeneity, nesting to respect production hierarchies, and deliberate fractionation to economize runs all collide in contemporary studies; missing plots and unequal replication often emerge as additional complications; see, for instance, Montgomery, (2020). At the center of every subsequent inference still lies a deceptively simple question: How many independent pieces of information—i.e., the degrees of freedom ($\mathrm{df}$)—does each factor really contribute? Once peripheral in balanced designs, $\mathrm{df}$ accounting now plays a central role in complex experimental analysis.
Classical rank additivity, first formalized by Cochran, (1941) for a single stratum of independent errors, states that for a total number $N$ of experimental observations, the (N1)(N-1) $\mathrm{df}$ split neatly among orthogonal treatment contrasts. However, the validity of Cochran’s identity is disrupted upon the introduction of a second randomization stage or the deliberate aliasing of higher-order interactions. Empirical practice tried to cope with these departures through moment-matching approximations suggested by Satterthwaite, (1946) and Kenward and Roger, (1997) or ad hoc sums-of-squares types devised by Searle et al., (1981). However, such remedies sacrifice both mathematical exactness and interpretability—often yielding non-integer, sometimes negative, $\mathrm{df}$ in heavily unbalanced layouts. Patterson and Williams, (1976) provide construction for lattices with complete and incomplete block designs, but $\mathrm{df}$ tables are only presented for balanced cases. Verbeke and Molenberghs, (2000) discuss how approximate $\mathrm{df}$ dominate certain design of experiments, despite these shortcomings.
Cornfield and Tukey, (1956) established the use of idempotent-projector calculus for ANOVA, concentrating on one-stratum models without aliasing; Hinkelmann and Kempthorne, (2005) present more recent applications of idempotent-projectors in modern experimental settings. A parallel tradition, begun by Yates, (1935) and systematized by Bailey, (2008), advocated idempotent projectors to represent randomization strata. While elegant for balanced split-plots, this projector calculus remained a craft discipline: each novel design demanded a fresh hand-drawn Hasse diagram (see, for example, (Bailey,, 2008, Chapter 4)), and unequal replication placed the method beyond reach.
Simultaneously, Box and Hunter, (1961) revolutionized $2^{k-p}$ factorial planning, but its celebrated resolution measure presupposed complete balance and a single source of error, and works only for regular, fully replicated 2–level fractions, without blocking. Pedagogical projector diagrams were presented by Bailey, (2008) to account for blocking; however, it lacks closed-form $\mathrm{df}$ under imbalance. Giesbrecht and Burns, (1985) compared exact versus Satterthwaite $F$-tests in split-plot large-sample asymptotics settings. The works of Searle et al., (1992); Harville, (1997); West et al., (2014) make advances in mixed-model effects, but all these treat $\mathrm{df}$ numerically rather than through analytic rank. The literatures on blocking, mixed models, and factorial aliasing have evolved in productive isolation. To the best of our knowledge, an exact, algebraic $\mathrm{df}$ identity that is simultaneously (i) multi-stratum, (ii) alias-aware, and (iii) robust to unequal replication has not been reported.
This paper reunifies those strands through a $\mathrm{df}$ partition theorem that operates at the intersection of algebra and group representation. For a set $\mathcal{S}$ that indexes all randomization strata in the experiment, let $\mathbf{P}_{s}$ denote the $N\times N$ idempotent projector that averages observations within randomization stratum $s\in\mathcal{S}$ and let $\mathcal{C}_{\overline{E}}$ denote the factorial-contrast linear subspace of $\mathbb{R}^{N}$ attached to the effect $\overline{E}$ (an alias class comprising all factorial effects that collapse into the same contrast space once the defining relations of a fractional design are imposed). By letting the family of stratum projectors $\mathbf{P}_{s}$} act jointly with the full level-permutation group $G$, we obtain a simultaneous diagonalization in which every fixed or random effect decomposes into orthogonal blocks $\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}$, where $\operatorname{Im}\mathbf{P}_{s}=\{\mathbf{P}_{s}\mathbf{v},\mathbf{v}\in\mathbb{R}^{N}\}$ is the column space of the projector $\mathbf{P}_{s}$. Essentially, when we intersect $\operatorname{Im}\mathbf{P}_{s}$ with a factorial-contrast space $\mathcal{C}_{\overline{E}}$, we isolate the part of effect $\overline{E}$ that is visible in stratum $s\in\mathcal{S}$. The dimension of this intersection $\sum\limits_{s}\sum\limits_{\overline{E}}\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$ is exactly the number $\mathrm{df}$ that the effect retains and can be tested against the mean square residing in that stratum. Crucially, the dimension of each block—and, therefore, its $\mathrm{df}$—can be written in closed form via orbit-stabilizer counts described by (Bailey,, 2008, Section 5.6), regardless of balance or aliasing.
The proposed theorem immediately produces the first exact $\mathrm{df}$ tables for the workhorses of mixed-model practice including the unbalanced split-plot, row-column, lattice, and general crossed-nested hierarchies. Where software has long fallen back on Satterthwaite or Kenward–Roger numbers, we supply integer ranks computable from small Gram matrices. These formulas both generalize and strictly contain the balanced textbook results, collapsing to them when replication is equal. Beyond enumeration, we introduce three quantitative aliasing metrics: The retention ratio $\rho(\overline{E})$ measures the proportion of an effect’s ideal $\mathrm{df}$ that survive blocking and fractionation; its complement $\delta(\overline{E})$ pinpoints the exact $\mathrm{df}$ deficiency, and the variance-inflation index $\alpha(\overline{E})\geq\rho(\overline{E})^{-1}$ bounds how severely any remaining contrasts are diluted. Taken together, they extend the Box–Hunter resolution (see Box and Hunter, (1961)) from regular two-level fractions to arbitrary level structures, stratum multiplicities, and missing data. Xu and Wu, (2001) analyze the evolution of resolution metrics for asymmetrical fractional factorial designs, but our rank-based work provides an algebraic alternative. Deng and Tang, (1999) extend Box–Hunter ideas for non-regular fractional designs and backs our claim that $\rho(\overline{E})$ generalizes resolution. Our projector–rank partition theorem including $\rho$, $\delta$ and $\alpha$ indices, therefore, fill a genuine methodological gap.
A trio of classical results emerges as corollaries from the projector–rank framework. (i) Single-stratum case: If the randomization consists of only one stratum $(\lvert\mathcal{S}\rvert=1,\mathbf{P}_{s_{1}}=\mathbf{I}_{N})$, the theorem reduces to Cochran’s rank-additivity identity $N-1=\sum\limits_{\overline{E}}\dim\mathcal{C}_{\overline{E}}$, reproducing exactly the fixed-effects result of Cochran, (1941). (ii) Balanced split-plot: Assigning factor $A$ to whole plots and factor $B$ to sub-plots with equal replication ($a$ levels, $b$ whole plots per level, $c$ sub-plots per whole plot) and substituting the corresponding projectors into the theorem yields the $\mathrm{df}$ table derived by Yates, (1935): (a1)(a-1) for $A$, a(b1)a(b-1) for whole-plot error, (c1)(c-1) for $B$, (a1)(c1)(a-1)(c-1) for $AB$, and a(b1)(c1)a(b-1)(c-1) for sub-plot error. (iii) Fractional-factorial resolution: For a regular $2^{k-p}$ fraction, the df-retention ratio satisfies $\rho(\overline{E})=0$ exactly when the effect order $|E|$ is at least the length of the shortest defining word. Hence, the Box–Hunter resolution $R$ of the design is $R=\min\left\{|E|:\rho(\overline{E})=0\right\}$, extending the classical definition to a rank-based statement that remains meaningful even when replication is unequal.
[…]
These theoretical advances translate into immediate practical benefits. Exact numerator and denominator $\mathrm{df}$ restore nominal type-I error control to $F$-tests in highly unbalanced agronomic and industrial trials. ANOVA diagnostic tables can flag unestimable contrasts before an experiment is run, preventing costly data-collection mistakes. Optimization heuristics can now search generator sets or blocking schemes by minimizing an integer-rank alias-loss index, avoiding computationally intensive Monte Carlo simulations. The work thus unifies disparate strands of design theory into a single algebraic framework and provides practitioners with precise $\mathrm{df}$—not approximations—for complex experiments.
The remainder of the paper is organized as follows. Section 2 states and proves the $\mathrm{df}$ partition theorem, developing the orbit-stabilizer dimension formula. Section 3 applies the theorem to derive exact $\mathrm{df}$ for the canonical mixed designs mentioned above. Section 3.4 formalizes the aliasing metrics $\rho$, $\delta$ and $\alpha$, and explores their connection to resolution and variance inflation. Section 4 presents three corollaries of the partition theorem. In Section 5, we illustrate the effectiveness of our method through empirical studies. Section 6 discusses some computational aspects together with concluding remarks.
2 $\mathrm{df}$ partition theorem
In practice we routinely introduce sequential randomization, mixed fixed and random effects, deliberate aliasing/confounding (fractional factorials, blocking on hard-to-change factors), and inevitable imbalance (missing plots, unequal replication). Under those departures the $\mathrm{df}$ tallies printed by software (Type I/II/III SS, Satterthwaite-Kenward approximations) are ad-hoc numerical expedients. They mask the underlying linear-algebraic fact: $\mathrm{df}$ are ranks of contrast subspaces intersected with stratum projectors. A precise theorem is developed in this section that (i) identifies those subspaces and (ii) adds their dimensions back to N1N-1, and therefore forms the foundation on which every subsequent $F$-test, variance-component estimator and information-loss metric must rest. The notation that will be used in what follows is summarized in Table 1.
Symbol Meaning
$k$ Number of factorial factors $F_{1},\dots,F_{k}$
$\Lambda_{i}$ Set of levels of $F_{i}$ with cardinality $\lvert l_{i}\rvert\geq 2$
$\Lambda=\Lambda_{1}\times\cdots\times\Lambda_{k}$ Full treatment grid
$\mathcal{S}$ Partially ordered set of randomization strata (Hasse diagram)
$\mathbf{Z}_{s}$ Indicator matrix mapping observations $\to$ units of stratum $s$
$\mathbf{P}_{s}=\mathbf{Z}_{s}(\mathbf{Z}_{s}^{\prime}\mathbf{Z}_{s})^{+}\mathbf{Z}_{s}^{\prime}$ Idempotent projector onto mean space of stratum $s$
$\bm{\Sigma}=\sum\limits_{s\in\mathcal{S}}\sigma_{s}^{2}\mathbf{P}_{s}$ Covariance of observational errors in a linear-mixed model
$\mathcal{A}\subseteq 2^{\{1,\dots,k\}}$ Alias ideal (defining words of a fractional design)
$\overline{\mathcal{E}}=2^{\{1,\dots,k\}}/\sim_{\mathcal{A}}$ Lattice of aliased factorial effects
$\tau:\overline{\mathcal{E}}\to\{\mathrm{fixed},\mathrm{random}\}$ Assignment of each aliased class to fixed/random side
Table 1: Notation for factorial design and mixed-model structure.
For an aliased effect class $\overline{E}\in\overline{\mathcal{E}}$, let $\mathcal{C}_{\overline{E}}=\{c\in\mathbb{R}^{\Lambda}:\sum_{\lambda\in\Lambda}c_{\lambda}=0,\sum_{\lambda:\lambda_{i}=\ell}c_{\lambda}=0\ \forall i\notin E\}$ denote the orthogonal complement w.r.t. the ordinary dot-product of lower-order marginal means; hence, it carries precisely the $\mathrm{df}$ of the factorial contrast $E$. The symmetric group $G=\prod\limits_{i}S_{l_{i}}$ acts (level-permutation) on $\Lambda$. The orbit-stabilizer theorem described by (Artin,, 1991, Proposition 6.8.4, pp. 179), (Rose,, 2009, Chapter 5) gives the dimension
$\dim(\mathcal{C}_{\overline{E}})$ $=$ $\frac{|G|}{|\operatorname{Stab}_{G}(\overline{E})|}\prod_{i\in E}(l_{i}-1),$ (1)
which collapses to $\prod\limits_{i\in E}(l_{i}-1)$ when $\mathcal{A}=\{\varnothing\}$, i.e., full factorial, no aliasing. Every projector $\mathbf{P}_{s}$ averages within the experimental units that share an identical label at stratum $s$ and eliminates contrasts that vary between such units. Hence, for a given factorial contrast $\mathcal{C}_{\overline{E}}$, the intersection $\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}$ is exactly the part of that contrast which survives into stratum $s$. Its dimension is the $\mathrm{df}$ we can test for $\overline{E}$ using the mean-square residing in $s$. Since $\bm{\Sigma}^{-1}=\sum\limits_{s}\sigma_{s}^{-2}\mathbf{P}_{s}$ is a linear combination of mutually orthogonal idempotents, it induces the generalized inner product $\langle\mathbf{u},\mathbf{v}\rangle_{\bm{\Sigma}}=\mathbf{u}^{\prime}\bm{\Sigma}^{-1}\mathbf{v}=\sum\limits_{s\in\mathcal{S}}\sigma_{s}^{-2}\mathbf{u}^{\prime}\mathbf{P}_{s}\mathbf{v}$. Under $\langle\cdot,\cdot\rangle_{\bm{\Sigma}}$ different strata are orthogonal ($\mathbf{P}_{s}\mathbf{P}_{s^{\prime}}=0$ for $s\neq s^{\prime}$), and treatment-label permutations commute with $\bm{\Sigma}^{-1}$. These two facts let us decompose any design matrix into orthogonal blocks simultaneously by stratum and by factorial effect. The $\mathrm{df}$ projector–rank partition theorem is as follows.
Theorem 2.1.
Let $\mathbf{X}_{\mathrm{fix}}=\bigoplus\limits_{\tau(\overline{E})=\text{fixed}}\!\mathbf{X}_{\overline{E}}$ and $\mathbf{Z}=\bigoplus\limits_{s\in\mathcal{S}}\mathbf{Z}_{s}$ be the fixed- and random-effects design matrices in a linear-mixed model whose treatment structure may be aliased and unbalanced. Then, with respect to the inner product $\langle\cdot,\cdot\rangle_{\bm{\Sigma}}$,
$\left.\begin{array}[]{rcl}\mathbf{X}_{\mathrm{fix}}&=&\bigoplus\limits_{s\in\mathcal{S}}\;\bigoplus\limits_{\begin{subarray}{c}\overline{E}\in\overline{\mathcal{E}}\\[1.0pt] \tau(\overline{E})=\mathrm{fixed}\end{subarray}}(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}),\\ \mathbf{Z}&=&\bigoplus\limits_{s\in\mathcal{S}}\;\bigoplus\limits_{\begin{subarray}{c}\overline{E}\in\overline{\mathcal{E}}\\[1.0pt] \tau(\overline{E})=\mathrm{random}\end{subarray}}(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}),\end{array}\right\}$ (4)
both of which are orthogonal direct sums. Therefore,
$N-1$ $=$ $\sum_{s\in\mathcal{S}}\sum_{\overline{E}\in\overline{\mathcal{E}}}\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right).$ (5)
Moreover, $\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)=\operatorname{rank}\left(\mathbf{X}_{\overline{E}}^{\!\prime}\mathbf{P}_{s}\mathbf{X}_{\overline{E}}\right)=\operatorname{trace}\left(\mathbf{P}_{\overline{E}}\mathbf{P}_{s}\right)$, where $\mathbf{P}_{\overline{E}}$ is the ordinary least-squares projector onto $\mathcal{C}_{\overline{E}}$.
Proof.
In the following, we remove the grand-mean column once from every matrix; this avoids carrying a “+1” term in each rank statement. Firstly, two observations share at most one highest stratum label, so the block-diagonal averaging performed by $\mathbf{P}_{s}$ and $\mathbf{P}_{s^{\prime}}$ falls on disjoint sets of rows. Therefore, $\mathbf{P}_{s}\mathbf{P}_{s^{\prime}}=0$ for $s\neq s^{\prime}$ iff strata are obtained by successive randomization, i.e., no observation belongs to two different units of the same level. Averaging step by step over every stratum amounts to identity because the final stratum is the set of individual observations. Thus, $\sum\limits_{s\in\mathcal{S}}\mathbf{P}_{s}=\mathbf{I}_{N}$. Since the projectors are pairwise orthogonal idempotents, their images form a direct (orthogonal) decomposition: $\mathbb{R}^{N}=\bigoplus\limits_{s\in\mathcal{S}}\operatorname{Im}\mathbf{P}_{s}$.
For each unordered effect $E\subseteq\{1,\dots,k\}$, the raw contrast space is defined by $\mathcal{C}_{E}=\{c\in\mathbb{R}^{\Lambda}:\sum_{\lambda}c_{\lambda}=0,\sum_{\lambda:\lambda_{i}=\ell}c_{\lambda}=0,\forall i\notin E\}$. Then, $\mathcal{C}_{E}\perp\mathcal{C}_{E^{\prime}}$ for $E\neq E^{\prime}$ and $\dim\mathcal{C}_{E}=\prod\limits_{i\in E}(l_{i}-1)$. This is because, a contrast for $E$ integrates to zero over levels of any factor not in $E$; inner products vanish because each term in one contrast is matched by a cancelling term in the other. Dimension is a direct-product count of linearly independent Helmert contrasts for each factor in $E$. If the design is fractionated with defining words $g_{1},\dots,g_{p}$, two effects $E_{1},E_{2}$ share identical columns in the design matrix precisely when $E_{1}\triangle E_{2}=(E_{1}\cup E_{2})\setminus(E_{1}\cap E_{2})$ is in the ideal generated by $\{g_{j}\}$. The quotient class is denoted $\overline{E}$. Any full-rank matrix $\mathbf{X}_{\overline{E}}$ whose columns span the aliased space is acceptable. For every $(\overline{E},s)$ the intersection $\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}$ is orthogonal to $\mathcal{C}_{\overline{E}^{\prime}}\cap\operatorname{Im}\mathbf{P}_{s^{\prime}}$ whenever $\overline{E}\neq\overline{E}^{\prime}$ or $s\neq s^{\prime}$. The orthogonality follows directly from the aforementioned paragraph. Furthermore, $\mathbf{P}_{s}\mathbf{P}_{s^{\prime}}=0$ implies $\operatorname{Im}\mathbf{P}_{s}\perp\operatorname{Im}\mathbf{P}_{s^{\prime}}$. Combined, the pairwise sums produce an orthogonal family of subspaces.
For any fixed $(\overline{E},s)$, $\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)=\operatorname{rank}\left(\mathbf{X}_{\overline{E}}^{\!\prime}\mathbf{P}_{s}\mathbf{X}_{\overline{E}}\right)$. This holds since the matrix $\mathbf{P}_{s}\mathbf{X}_{\overline{E}}$ has image contained in $\operatorname{Im}\mathbf{P}_{s}$. Its column-space is exactly the intersection because any column of $\mathbf{X}_{\overline{E}}$ lies in $\mathcal{C}_{\overline{E}}$, and applying $\mathbf{P}_{s}$ projects it to the largest piece living inside the stratum image. The rank of a matrix equals the dimension of its column space, giving the result. Deleting the grand mean, the orthogonal direct sum spans $\mathbb{R}^{N-1}$. Consequently, $N-1=\sum\limits_{s,E}\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$. This is valid since orthogonality of the blocks implies that dimensions add. Nothing in $\mathbb{R}^{N-1}$ is missing because (i) each observation can be written as its successive deviations from higher to lower strata, and (ii) every deviation decomposes uniquely into factorial contrasts. Therefore, the identity $N-1=\sum\limits_{s,E}\operatorname{rank}\left(\mathbf{X}_{\overline{E}}^{\!\prime}\mathbf{P}_{s}\mathbf{X}_{\overline{E}}\right)$ follows. This completes the proof of Theorem 2.1. ∎
Note that, $\mathcal{C}_{\overline{E}}$ is obtained from the orbit-stabilizer or Kronecker product formula. $\operatorname{Im}\mathbf{P}_{s}$ denotes the subspace of responses visible at stratum $s$ and can be computed by group-wise averaging. $\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}$ is the portion of $\overline{E}$ that survives into stratum $s$ and is given by the row/column rank of $\mathbf{X}_{\overline{E}}^{\!\prime}\mathbf{P}_{s}$. $\dim(\cdot)$ is the $\mathrm{df}$ assigned to effect $\overline{E}$ and tested against mean square for stratum $s$, and can be computed using any numerical rank routine; closed forms are given later. Since (5) is an identity, every downstream test (fixed-effect $F$-test, variance ratio, likelihood-ratio) inherits exact numerator and denominator $\mathrm{df}$. If $\sum\limits_{E,s}\dim(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s})<N-1$, the deficit pinpoints sources of aliasing loss or singular randomization, thereby aiding model diagnostics. Competing designs (different fractional generators, blocking plans) can be scored by how their $\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$ tallies distribute $\mathrm{df}$ across strata-an information-theoretic design-optimality criterion, thus facilitating design comparison. Theorem 2.1 provides a unified framework—subsuming Cochran’s one-stratum theorem, Yates’ Hasse diagram, and Box–Hunter resolutions—while (4) provides a single algebraic object on which extensions (split-plot, nested, strip-plot, repeated measures) are mere changes of the projector family $\{\mathbf{P}_{s}\}$.
Substituting explicit $\mathbf{P}_{s}$ and $\mathbf{X}_{\overline{E}}$ into $\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$ yields closed-form $\mathrm{df}$ for split-plot, Latin-square, crossed-nested mixed models. The ratio of $\sum\limits_{s}\dim(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s})$ to $\dim(\mathcal{C}_{\overline{E}})$ provides the aliasing-loss ratio $\rho(\overline{E})$. Block-orthogonality of (4) simplifies likelihood gradients and results in exact restricted or residual maximum likelihood (REML variance-component score equations. Generalized $F$-tests with exact $\mathrm{df}$ without Satterthwaite approximation can be obtained by setting the numerator to be the rank of $\mathbf{X}^{\prime}\mathbf{P}_{s}\mathbf{X}$ and the denominator to be the rank deficit of residual space. Optimal fractionation/confounding search algorithms can be devised by maximizing $\sum\limits_{E}w_{E}\rho(\overline{E})$ under run budget, with $\sum\limits_{E}w_{E}=1$ where $w_{E}$ is simply a user-supplied weight expressing how much the user cares about preserving information for a particular factorial effect $E$. In summary, once the orthogonal decomposition (4) is established, exact tests, information-loss metrics, and design search heuristics can be derived systematically using algebraic identities rather than ad-hoc approximation.
3 Exact $\mathrm{df}$ formulas for classical mixed designs
In Section 2, we saw that every experimental layout can be analyzed by intersecting two linear objects, namely, factorial-contrast space $\mathcal{C}_{\overline{E}}$ and stratum projector $\mathbf{P}_{s}$, and that $\text{$\mathrm{df}$}(\overline{E},s)=\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)=\operatorname{rank}\!\left(\mathbf{X}_{\overline{E}}^{\!\prime}\mathbf{P}_{s}\mathbf{X}_{\overline{E}}\right)$. In this section, we put this identity to work for three families of designs that cover mixed-model practices, such as split-plot and its variants (split-split, strip-plot), row-column structures (Latin squares, lattice, incomplete-block) and crossed-nested hierarchies (animal models, multi-site trials, education data); see, for example, Milliken and Johnson, (2009). We derive closed-form ranks-no iterative numerical approximations, no Satterthwaite or Kenward–Roger guesses are required. These formulas give exact $F$-test numerators/denominators, guaranteeing the advertised type I error; they feed directly into REML/ML information matrices, yielding analytic standard-error formulas; and they reveal how unequal replication or missing plots erode information, thereby providing a design-diagnostics tool.
3.1 Split-plot (one whole-plot factor $A$, one sub-plot factor $B$)
Source Contrast basis df
AA (whole-plot) $\mathcal{C}_{A}\otimes\mathbf{1}_{b}\otimes\mathbf{1}_{c}$ a1a-1
Whole-plot error $\mathbf{1}_{a}\otimes\mathcal{C}_{b(\text{plots})}\otimes\mathbf{1}_{c}$ a(b1)a(b-1)
$B$ (sub-plot) $\mathbf{1}_{a}\otimes\mathbf{1}_{b}\otimes\mathcal{C}_{B}$ c1c-1
ABAB $\mathcal{C}_{A}\otimes\mathbf{1}_{b}\otimes\mathcal{C}_{B}$ (a1)(c1)(a-1)(c-1)
Sub-plot error Orthogonal complement abc(b1)abc(b-1)
Table 2: $\mathrm{df}$ and contrast bases in a split-plot design.
First, consider the balanced case. The whole plots comprise $a$ levels of $A$; each level replicated in $b$ whole plots. For the sub-plots, every whole plot is split into $c$ sub-plots receiving $B$ levels exactly once. The strata comprises the following: $\mathcal{S}=\{s_{\text{whole}},\;s_{\text{sub}}\}$, $\mathbf{P}_{\text{whole}}=\operatorname{diag}\left(\tfrac{1}{bc}\mathbf{1}_{bc}\right)_{\!a}$, and $\mathbf{P}_{\text{sub}}=\mathbf{I}-\mathbf{P}_{\text{whole}}$. The factor-induced contrast spaces are Kronecker products of one-way contrasts with all-1 vectors. The implications of orthogonality on the $\mathrm{df}$ and contrast bases in a split-plot design are summarized in Table 2. Plugging the bases into $\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)$ yields the $\mathrm{df}$. Since the design is balanced, each rank is the product of the component dimensions.
For the case of unequal replication ($n_{ij}$ sub-plots in whole plot $i$, level $j$ of $B$), let $\mathbf{R}_{A}=\left[\bar{y}_{i\cdot\cdot}-\bar{y}_{\cdot\cdot\cdot}\right]_{1\times a}$, $\mathbf{R}_{B}=\left[\bar{y}_{\cdot\cdot j}-\bar{y}_{\cdot\cdot\cdot}\right]_{1\times c}$, and $\mathbf{R}_{AB}=\left[\bar{y}_{ij\cdot}-\bar{y}_{i\cdot\cdot}-\bar{y}_{\cdot\cdot j}+\bar{y}_{\cdot\cdot\cdot}\right]_{a\times c}$. Since $\mathbf{P}_{\text{whole}}$ averages within whole plots, $\mathbf{X}_{\!A}^{\prime}\mathbf{P}_{\text{whole}}\mathbf{X}_{\!A}=\mathbf{R}_{A}\mathbf{R}_{A}^{\prime}$, and $\operatorname{rank}(\mathbf{R}_{A})=\lvert\{i:\bar{y}_{i\cdot\cdot}\neq\bar{y}_{\cdot\cdot\cdot}\}\rvert$. The other ranks follow analogously, thereby giving the $\mathrm{df}$. The key point is that even with grotesque imbalance the $\mathrm{df}$ are still integers computable by ordinary rank on small summary matrices, with no heuristics involved.
3.2 Row-column (Latin square and variants)
For the balanced Latin square ($r=c=t$, no missing cells), only one stratum exists, so $\mathbf{P}=\mathbf{I}$. The design matrix partitions as $\mathbf{X}=\left[\mathbf{X}_{\!R}\mid\mathbf{X}_{\!C}\mid\mathbf{X}_{\!T}\right]$, $\operatorname{rank}(\mathbf{X})=N-1=r^{2}-1$. Orthogonality yields the $\mathrm{df}$: $\text{Rows}=r-1$, $\text{Columns}=c-1$, $\text{Treatments}=t-1$, and $\text{Error}=(r-1)(c-2)$. For the case of missing cells or unequal replication, write $\mathbf{M}$ for the $r\times c$ incidence matrix whose $(i,j)$ entry is the number of observations in row $i$, column $j$. Pukelsheim, (2006) showed that, since each structural zero or replication imbalance removes one independent contrast, the $\operatorname{nullity}\left(\mathbf{M}-\tfrac{1}{r}\mathbf{1}_{r}\mathbf{1}_{r}^{\prime}\mathbf{M}\right)$, $\operatorname{nullity}\left(\mathbf{M}-\mathbf{M}\tfrac{1}{c}\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}\right)$ and $\operatorname{nullity}(\mathbf{M})-\operatorname{nullity}\left(\mathbf{M}-\tfrac{1}{r}\mathbf{1}_{r}\mathbf{1}_{r}^{\prime}\mathbf{M}\right)-\operatorname{nullity}\left(\mathbf{M}-\mathbf{M}\tfrac{1}{c}\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}\right)$ counts the lost $\mathrm{df}$ exactly. Later we will see that these lost $\mathrm{df}$, in fact, constitute the $\mathrm{df}$-deficiency.
3.3 General crossed-nested mixed models
Let $\mathcal{N}$ and $\mathcal{C}$ denote the nesting tree of random factors and the lattice of crossed fixed factors, respectively. Define for every random stratum $s\in\mathcal{N}$ its projector $\mathbf{P}_{s}$. For every factorial contrast $\overline{E}$, let $\mathbf{X}_{\overline{E}}$ be the usual dummy-coded matrix.
Proposition 1.
If the design is orthogonally generated, i.e., each fixed factor is crossed with every ancestor in the nesting tree, and the replication within each terminal unit is constant, then $\mathrm{df}(s,E)=\prod\limits_{i\in E}(l_{i}-1)\prod\limits_{\begin{subarray}{c}r\prec s\\ r\in\mathcal{N}\end{subarray}}(n_{r}-1)$, where $n_{r}$ is the number of children per unit at level $r$.
Proof.
Both $\mathbf{P}_{s}$ and $\mathbf{X}_{\overline{E}}$ are Kronecker products of 11-way projectors. The rank of their Gram product is the product of the ranks of each Kronecker factor. ∎
If replication is unequal, replace the integer factors $n_{r}-1$ by the rank of the corresponding incidence matrix inside stratum $s$. Note that, $\mathrm{df}(s,E)$ is the exact number of $\mathrm{df}$ that the factorial effect $E$ retains in stratum $s$. Here, $E$ is an un-aliased factorial effect, i.e., a subset of the factor indices $\{1,\dots,k\}$ and $l_{i}$ is the number of levels of factor $F_{i}$. The factor-contrast space for $F_{i}$ has dimension $l_{i}-1$, while $s$ is a node in the nesting tree $\mathcal{N}$ representing one randomization stratum. $r\prec s$ denotes the nodes $r$ that are strict ancestors of $s$ in the nesting tree, i.e., random factors that contain stratum $s$. $n_{r}$ is the constant number of sub-units per unit at stratum $r$.
The treatment-contrast component $\prod_{i\in E}(l_{i}-1)$ gives the dimension of the pure factorial contrast space for $E$ in a full, unblocked design. For example, if $E=\{A,C\}$ with $l_{A}=3$ and $l_{C}=4$ levels, the factor part contributes (31)(41)=6(3-1)(4-1)=6 $\mathrm{df}$. The replication (stratum) component is $\prod_{r\prec s}(n_{r}-1)$. Each ancestor level $r$ in the nesting tree introduces $n_{r}-1$ independent directions within the units of $r$. Multiplying these factors counts how many “within-unit” contrasts survive down to stratum $s$. For example, in an animal model Sire $\to$ Dam $\to$ Animal (see the book by Henderson, (1984) for worked animal-model examples.), the $\mathrm{df}$ for a diet effect observed at the Animal stratum multiplies $(n_{\text{Sire}}-1)(n_{\text{Dam}}-1)$. For every pair $(s,E)$ the value $\mathrm{df}(s,E)$ feeds into the global identity $N-1=\sum_{s\in\mathcal{S}}\sum_{E}\mathrm{df}(s,E)$, so that the numerator $\mathrm{df}$ for the $F$-test of $E$ against stratum $s$ is precisely $\mathrm{df}(s,E)$. The residual $\mathrm{df}$ for $s$ are obtained by subtracting all $\mathrm{df}(s,E)$ allocated to that stratum from its total dimension. Hence $\mathrm{df}(s,E)$ is the building block that turns the abstract projector–rank theorem into explicit, integer $\mathrm{df}$ for each source in the ANOVA table.
3.4 Quantifying aliasing loss
In this section, we quantify how much of each factorial contrast survives after blocking, fractionation, or imbalance. We begin with a quick recap of aliasing. There are three broad types of aliasing: (a) In a fractional factorial the defining relation $\mathcal{A}=\left\langle\;ABCD,\;AE,\;\dots\right\rangle$ identifies factor-subset pairs whose contrast spaces coincide, i.e., $E_{1}\sim_{\mathcal{A}}E_{2}\;\Longleftrightarrow\;E_{1}\triangle E_{2}\in\mathcal{A}$. Orthogonality of the full $2^{k}$ lattice collapses to a quotient lattice $\overline{\mathcal{E}}=2^{\{1,\dots,k\}}/\!\sim_{\mathcal{A}}$. This is called structural aliasing; see, Box and Hunter, (1961). (b) In randomization aliasing (see, Yates, (1935)), each projector $\mathbf{P}_{s}$ averages observations within the units of stratum $s$. Any contrast lying in $\ker(\mathbf{P}_{s})$ is filtered out at that stratum; if every stratum eliminates it, the effect is unestimable even in the full factorial. (c) In imbalance-induced aliasing, missing plots or unequal replication destroy orthogonality, i.e., the Gram matrix $\mathbf{X}_{\overline{E}}^{\prime}\mathbf{X}_{\overline{E}}$ loses rank, merging some contrasts into others. Hence, aliasing loss is multi-dimensional: $\mathrm{df}$ loss measures how many independent directions disappear; information loss quantifies how much the variance of the remaining directions inflates; and bias indicates how strongly an unmeasurable effect contaminates estimable ones.
From Theorem 2.1, $\mathrm{df}(s,E)=\dim\left(\mathcal{C}_{\overline{E}}\cap\operatorname{Im}\mathbf{P}_{s}\right)=\operatorname{rank}\left(\mathbf{X}_{\overline{E}}^{\prime}\mathbf{P}_{s}\mathbf{X}_{\overline{E}}\right)$. The total ideal $\mathrm{df}$ is $d_{E}^{\text{ideal}}=\dim(\mathcal{C}_{\overline{E}})$ (e.g., for a two-level factor $E$ it is $1$; for a three-level factor it is $2$, etc.). The recoverable $\mathrm{df}$ is $d_{E}^{\text{obs}}=\sum_{s}\mathrm{df}(s,E)$. The information-retention ratio is defined by $\rho(E)=\frac{d_{E}^{\text{obs}}}{d_{E}^{\text{ideal}}}\in[0,1]$ and the $\mathrm{df}$-deficiency is defined by $\delta(E)=d_{E}^{\text{ideal}}-d_{E}^{\text{obs}}=\dim\left(\ker\Pi_{E}\right)$, $\Pi_{E}=\sum\limits_{s}\mathbf{P}_{s}\mathbf{X}_{\overline{E}}$. It can be seen that $\rho(E)=1$ iff the effect is fully estimable (no $\mathrm{df}$ lost), $\rho(E)=0$ iff $E$ is completely aliased or eliminated, and $0<\rho(E)<1$ indicates partial aliasing-typical when we have both fractionation and missing plots. Since $\sum\limits_{E}\delta(E)$ equals the global shortfall in N1N-1 $\mathrm{df}$, the set $\{\delta(E)\}_{E}$ pinpoints which effects paid the price.
$\mathrm{df}$ say how many directions survive; to measure quality we look at the alias (or dispersion) matrix $\mathbf{A}(E)=\left(\mathbf{X}_{\overline{E}}^{\prime}\bm{\Sigma}^{-1}\mathbf{X}_{\overline{E}}\right)^{+}\mathbf{X}_{\overline{E}}^{\prime}\bm{\Sigma}^{-1}(I-\sum_{s}\mathbf{P}_{s})\bm{\Sigma}^{-1}\mathbf{X}_{\overline{E}}$, where $\mathbf{A}(E)=0$ when $E$ is orthogonal to every nuisance and random effect and nonzero eigenvalues of $\mathbf{A}(E)$ inflate the variance of $\hat{\beta}_{E}$ or, if $E$ is unestimable, represent aliasing bias projected into other effects. A scalar aliasing index
$\alpha(E)$ $=$ $\frac{\operatorname{trace}\mathbf{A}(E)}{\operatorname{trace}\left(\left(\mathbf{X}_{\overline{E}}^{\prime}\bm{\Sigma}^{-1}\mathbf{X}_{\overline{E}}\right)^{+}\right)}\in[0,\infty)$
acts as a variance-inflation factor (VIF) for the whole contrast space. When $\rho(E)<1$ we always have $\alpha(E)=\infty$ on the lost directions; otherwise $\alpha(E)\geq 1$.
The information-retention ratio $\rho(E)$ and the $\mathrm{df}$-deficiency $\delta(E)$ extend classical ideas built for balanced, regular designs to any mixture of fractionation, blocking, and imbalance. For instance, consider the Box–Hunter resolution $R$. For regular $2^{k-p}$ fractions, $R$ is the length of the shortest defining word. The rule “a factor of length $q$ is aliased at most with interactions of order RqR-q” translates exactly to $\rho(E)=2^{-(R-|E|)}$ (balanced, regular fraction). Similarly, the generalized resolution proposed by Cheng and Ye, (2004) measures the minimax cosine between contrast subspaces; $\rho(E)$ is the rank analogue of the same idea and reduces to their criterion when each stratum is a full average over generators. Minimizing total $\mathrm{df}$ deficiency $\sum\limits_{E}w_{E}\delta(E)$ with suitable weights $w_{E}$ reproduces minimum aberration search heuristics used for design construction.
3.5 An example
Consider a design with $2^{3-1}$ fraction defined by $ABC=+1$ having factors $A,B,C$ and 8 runs. There are two blocks of 4 runs, with blocks confound factor $B$. Each main effect has 11 $\mathrm{df}$, and the two-factor interactions have 11 $\mathrm{df}$ each. The three-factor interaction is aliased with mean. $\mathbf{P}_{\text{block}}$ averages runs within each 4-run block. The retention is as follows: Factor $A$ has contrast vector orthogonal to blocks, thus $\rho(A)=1$. Factor $B$ is fully confounded leading to $\mathrm{df}(\text{block},B)=0$, so $\rho(B)=0$, while $AB$ inherits the deficiency of $B$ because $ABC=+1$ which implies $\rho(AB)=0$. The index table is shown in Table 3, where we see that the total $\mathrm{df}$ lost =3=3, matching the block + generator restrictions.
Effect $\bm{\rho}$ $\bm{\delta}$ Interpretation
AA 1 0 fully estimable
$B$ 0 1 lost to blocking
$C$ 1 0 estimable
ABAB 0 1 lost via $B$ + defining relation
ACAC 1 0 estimable
BCBC 0 1 lost
Mean, ABCABC - - absorbed in intercept
Table 3: Effect estimability under aliasing and blocking.
3.6 Why aliasing-loss metrics are essential
[…]
The eigenstructure of $\mathbf{A}(E)$ plugs into closed-form REML standard-error formulas; contrasts with $\rho(E)\ll 1$ dominate the uncertainty budget. A single scalar criterion referred to as alias-loss index (ALI), given by $\text{ALI}(D)=\sum_{E}w_{E}\left(1-\rho(E)\right)$ ranks candidate designs; combinatorial search (genetic algorithms, IP solvers) can minimize ALI subject to run limits and blocking rules, thereby aiding optimal design under constrained budgets. These metrics could provide software diagnostics, by seamlessly integrating $\rho(E)$, $\delta(E)$, $A(E)$ and $\alpha(E)$ in statistical packages (e.g., FrF2, AlgDesign, doebioresearch). With the aliasing-loss machinery in place, exact variance-component estimation, power analysis, and algorithmic design optimisation reduces to substituting $\rho,\delta,$ or $\alpha$ into familiar mixed-model formulas.
4 Corollaries
The following three well-known results emerges as corollaries from the projector–rank framework: Cochran’s identity, $\mathrm{df}$ for Yates’ balanced split-plot and Box–Hunter resolution.
Corollary 1.
Cochran’s identity: If the design has a single randomization stratum and is fully balanced and orthogonal, then Theorem 2.1 reduces to identity $N-1=\sum_{E}\dim\mathcal{C}_{E}$.
Proof.
For a single stratum with $\lvert\mathcal{S}\rvert=1$, the projector family reduces to $\mathbf{P}_{s_{0}}=\mathbf{I}_{N}$. Therefore, for every effect $E$, $\dim\left(\mathcal{C}_{E}\cap\operatorname{Im}\mathbf{P}_{s_{0}}\right)=\dim\left(\mathcal{C}_{E}\cap\mathbf{I}_{N}\right)=\dim\mathcal{C}_{E}$. Since the design is balanced and complete factorial, the contrast spaces $\{\mathcal{C}_{E}\}$ are pairwise orthogonal and $\bigoplus\limits_{E}\mathcal{C}_{E}=\left(\mathbf{1}_{N}\right)^{\!\perp}\subset\mathbb{R}^{N}$, whose dimension is N1N-1. Substituting these results into Theorem 2.1 yields Cochran’s identity: $N-1=\sum\limits_{E}\dim\mathcal{C}_{E}$. This completes the proof of Corollary 1. ∎
Corollary 2.
Yates’ balanced split-plot $\mathrm{df}$: Let an experiment have a whole-plot (fixed) factor $A$ with $a\geq 2$ levels, $b\geq 2$ whole-plot replicates per level of $A$, and a sub-plot (fixed) factor $B$ with $c\geq 2$ levels applied once in each whole plot, so that the total number of observations is $N=abc$. Denote the two randomization strata by $s_{\mathrm{w}}\;(\text{whole plots})$ and $s_{\mathrm{s}}\;(\text{sub-plots})$, with idempotent projectors $\mathbf{P}_{\mathrm{w}}$ and $\mathbf{P}_{\mathrm{s}}=\mathbf{I}-\mathbf{P}_{\mathrm{w}}$. Let $\mathcal{C}_{A},\mathcal{C}_{B},\mathcal{C}_{AB}$ be the usual balanced-contrast subspaces for the effects $A$, $B$, and $AB$, respectively. Then, as a direct consequence of Theorem 2.1,
$\dim\left(\mathcal{C}_{E}\cap\operatorname{Im}\mathbf{P}_{s}\right)=\begin{cases}a-1,(E,s)=(A,s_{\mathrm{w}}),\\ c-1,(E,s)=(B,s_{\mathrm{s}}),\\ (a-1)(c-1),(E,s)=(AB,s_{\mathrm{s}}),\\ 0,\text{otherwise}.\end{cases}$
Hence the balanced split-plot ANOVA table is given by Table 4.
Source $\mathrm{df}$
Whole-plot factor AA a1a-1
Whole-plot error a(b1)a(b-1)
Sub-plot factor $B$ c1c-1
ABAB interaction (a1)(c1)(a-1)(c-1)
Sub-plot error a(b1)(c1)a(b-1)(c-1)
Total abc1abc-1
Table 4: $\mathrm{df}$ for a split-plot design.
This reproduces the allocation given by (Yates,, 1935, Section 5) for balanced split-plots, confirming that the classical result is the one-stratum-per-effect special case of the general projector–rank framework.
Proof.
The randomization strata are the whole-plot stratum $s_{\mathrm{w}}$ with units $(i,j)$ of size $c$ and the sub-plot stratum $s_{\mathrm{s}}$ with individual observations $(i,j,k)$. Define $\mathbf{P}_{\mathrm{w}}=I_{a}\;\otimes\;I_{b}\;\otimes\;\tfrac{1}{c}\mathbf{J}_{c}$ and $\mathbf{P}_{\mathrm{s}}=I_{abc}-\mathbf{P}_{\mathrm{w}}$, where $\mathbf{J}_{c}$ is the $c\times c$ all-ones matrix. Let $\mathcal{J}_{n}=\operatorname{span}\{\mathbf{J}_{n}\mathbf{1}_{n}\}$, $\mathcal{C}_{n}=\mathcal{\mathbf{missing}}{J}_{n}^{\perp}$, $\dim\mathcal{C}_{n}=n-1$. Then, the basis for the effects $A$, $B$ and $AB$ are $\mathcal{C}_{a}\otimes\mathcal{\mathbf{missing}}{\mathbf{J}}_{b}\otimes\mathcal{\mathbf{missing}}{\mathbf{J}}_{c}$ with dimension a1a-1, $\mathcal{\mathbf{missing}}{\mathbf{J}}_{a}\otimes\mathcal{\mathbf{missing}}{\mathbf{J}}_{b}\otimes\mathcal{C}_{c}$ with dimension c1c-1 and $\mathcal{C}_{a}\otimes\mathcal{\mathbf{missing}}{\mathbf{J}}_{b}\otimes\mathcal{C}_{c}$ with dimension (a1)c(c1)(a-1)c(c-1), respectively. Write $\mathbf{X}_{A},\;\mathbf{X}_{B},\;\mathbf{X}_{AB}$ for any full-rank matrices whose columns span these subspaces. A direct consequence of Theorem 2.1, we have the following.
First, consider the whole-plot stratum $s_{\mathrm{w}}$. Since $\mathbf{P}_{\mathrm{w}}\left(\mathcal{\mathbf{missing}}{\mathbf{J}}_{b}\otimes\mathcal{\mathbf{missing}}{\mathbf{J}}_{c}\right)=\mathcal{\mathbf{missing}}{\mathbf{J}}_{b}\otimes\mathcal{\mathbf{missing}}{\mathbf{J}}_{c}$, $\operatorname{Im}\mathbf{P}_{\mathrm{w}}=\mathcal{\mathbf{missing}}{\mathbf{J}}_{a}\otimes\mathcal{\mathbf{missing}}{\mathbf{J}}_{b}\otimes\mathcal{\mathbf{missing}}{\mathbf{J}}_{c}$. For factor $A$, $\mathbf{X}_{A}$ is constant within whole plots. Therefore, $\mathbf{X}_{A}^{\!\prime}\mathbf{P}_{\mathrm{w}}\mathbf{X}_{A}=\mathbf{X}_{A}^{\!\prime}\mathbf{X}_{A}$ (full rank), which implies $\mathrm{df}$$(A,s_{\mathrm{w}})=a-1$. Factors $B$ and $AB$ each varies within whole plots. Therefore, $\mathbf{P}_{\mathrm{w}}\mathbf{X}_{B}=\mathbf{P}_{\mathrm{w}}\mathbf{X}_{AB}=0$ which implies no $\mathrm{df}$ at $s_{\mathrm{w}}$. With the number of whole plots =ab=ab, removing the grand mean and $A$ leaves $\mathrm{df}$$(\text{W.Err})=ab-1-(a-1)=a(b-1)$ for the whole-plot error. Next, consider the sub-plot stratum $s_{\mathrm{s}}$. Here, $\mathbf{P}_{\mathrm{s}}=\mathbf{I}-\mathbf{P}_{\mathrm{w}}$. For factor $A$, $\mathbf{P}_{\mathrm{s}}\mathbf{X}_{A}=0$ which implies no $\mathrm{df}$. For factor $B$, $\mathbf{P}_{\mathrm{s}}\mathbf{X}_{B}=\mathbf{X}_{B}$ (already orthogonal to whole-plot means), which implies $\mathrm{df}$ $(B,s_{\mathrm{s}})=c-1$. For factor $AB$, similarly $\mathrm{df}$$(AB,s_{\mathrm{s}})=(a-1)(c-1)$. The residual $\mathrm{df}$ in this stratum are $\mathrm{df}$ $(\text{S.Err})=N-1-[(a-1)+(a(b-1))+(c-1)+(a-1)(c-1)]=a(b-1)(c-1)$ for the sub-plot error. These results exactly match the $\mathrm{df}$ for a split-plot design summarized in Table 4. This completes the proof of Corollary 2. ∎
Corollary 3.
Box–Hunter resolution: Let the experiment be a regular $2^{k-p}$ fractional factorial with independent defining words $g_{1},\dots,g_{p}\in 2^{\{1,\dots,k\}}$; the Box and Hunter, (1961) resolution $R$ be $R=\min_{j}|g_{j}|$, the length of the shortest defining word; there be one randomization stratum ($|\mathcal{S}|=1,\;\mathbf{P}_{s_{0}}=\mathbf{I}_{N}$); and $\rho(\overline{E})=\dfrac{\dim\left(\mathcal{C}_{\overline{E}}\cap\mathbf{I}_{N}\right)}{\dim\mathcal{C}_{\overline{E}}}$ be the $\mathrm{df}$ retention ratio. Then $R=\min\left\{|E|\;:\;\rho(\overline{E})=0\right\}$.
Proof.
In a regular fraction the alias ideal generated by the defining words is given by $\mathcal{A}=\langle g_{1},\dots,g_{p}\rangle\subset 2^{\{1,\dots,k\}}$. For any factorial effect $E$ the aliased class is $\overline{E}=E+\mathcal{A}$. Since the design is balanced and two-level, the contrast space for a single effect has dimension $\dim\mathcal{C}_{E}=1$. With $\mathbf{P}_{s_{0}}=\mathbf{I}_{N}$, the numerator of $\rho$ is $\dim(\mathcal{C}_{\overline{E}})$ if $\mathcal{C}_{\overline{E}}\neq\{0\}$ and zero otherwise. Hence, $\rho(\overline{E})=1$ if $E\notin\mathcal{A}$ and $\rho(\overline{E})=0$ if $E\in\mathcal{A}$. An effect (alias class) has $\rho=0$ iff the effect itself lies in the ideal: $\rho(\overline{E})=0\Longleftrightarrow E\in\mathcal{A}\Longleftrightarrow\exists g_{j}$ such that $E=g_{j}$ or $\lvert E\rvert\geq\lvert g_{j}\rvert$ for some $j$. The shortest effect contained in $\mathcal{A}$ is precisely the shortest defining word, whose length by definition equals $R$. Therefore $\min\left\{\lvert E\rvert:\rho(\overline{E})=0\right\}=R$. The resolution-$R$ criterion (“length of the shortest effect aliased with the mean”) is identical to the statement “the smallest effect order whose $\mathrm{df}$-retention ratio vanishes.” Thus, resolution is obtained as an immediate corollary of the $\mathrm{df}$-retention metric. This completes the proof of Corollary 3. ∎
5 Experimental studies
We created a canonical split-plot experiment with a whole-plot factor $A$ (3 levels, $a=3$) and a sub-plot factor $B$ (4 levels, $c=4$). Each level of $A$ contained $b=4$ whole plots, giving $N=abc=48$ potential observations in the balanced layout. Random variability is introduced exactly as mixed-model textbooks prescribe: a whole-plot (WP) effect sampled i.i.d. $N(0,\sigma_{\text{WP}}^{2})$ with $\sigma_{\text{WP}}=1$, and an independent residual $N(0,\sigma_{\text{res}}^{2})$ with $\sigma_{\text{res}}=1$. To mimic imbalance we randomly delete 20% of the sub-plots. Since no fixed effects were added to the data-generation equation, the global null $H_{0}\colon A$ has no effect is true by construction. Practitioners routinely test $A$ by forming $F=\frac{\mathrm{MS}_{A}}{\mathrm{MS}_{\text{sub-plot}}}$ and quoting denominator $\mathrm{df}$ from a completely randomized factorial (the Type I or III ANOVA default in many packages). Our projector–rank framework instead assigns the $A$ numerator to the WP stratum and uses the whole-plot mean square as denominator, with its exact $\mathrm{df}$ $\nu\_2=a(b-1)=12$. Hence, we compared (i) Exact $\mathrm{df}$ test—denominator $\mathrm{df}=12$ balanced, $12$ after deletions (integer, stratum-correct) and (ii) Naive test—denominator $\mathrm{df}$ = residual $\mathrm{df}$ from a two-way ANOVA ($\approx 36$ balanced, $\approx 27$ with 20% missing). Each version was applied to 1000 Monte Carlo replicates at nominal $\alpha=0.05$.
Results are tabulated in Table 5.
Scenario
Exact $\mathrm{df}$
empirical size
Naïve
empirical size
Balanced split-plot 0.040 0.529
20% missing sub-plots 0.051 0.466
Table 5: Empirical size using exact $\mathrm{df}$ versus the naïve approach under two scenarios.
With the projector–rank $\mathrm{df}$, the empirical type-I error closely matches the nominal 5% target in both layouts (4.0% and 5.1%). Using the naïve denominator $\mathrm{df}$ multiplies the false positive rate roughly ten-fold, reaching 52.9% even when the design is perfectly balanced and remaining disastrously high (46.6%) after imbalance is introduced. The split-plot hierarchy places all whole-plot variation in a higher stratum than sub-plot error. Treating the entire design as if every observation were independent ignores this structure, inflates the denominator $\mathrm{df}$, and produces severely liberal $F$-tests. The $\mathrm{df}$ partition theorem rectifies the problem automatically: by intersecting the AA-contrast space with the whole-plot projector $P_{\text{WP}}$ it delivers the correct integer $\mathrm{df}$ regardless of missing cells. This simulation, therefore, illustrates the practical payoff of the exact $\mathrm{df}$: researchers avoid spurious discoveries that would be almost certain under standard approximate analyses.
Next, we consider the nested animal model (see, for example, Henderson, (1984)). Animal-breeding studies often follow a three-tier hierarchy: sires at the top, dams nested within sires, and individual animals within dams. To mirror that structure, we simulated 6 sires, each mated to 4 dams, with each dam producing 3 offspring, for a potential total of $6\times 4\times 3=72$ observations. Random effects were generated at both upper levels: sire effects $\mathcal{N}(0,\sigma^{2}_{\text{sire}})$ with $\sigma_{\text{sire}}=1$, and dam-within-sire effects $\mathcal{N}(0,\sigma^{2}_{\text{dam}})$ with $\sigma_{\text{dam}}=0.5$. Residual variation at the animal level followed $\mathcal{N}(0,\sigma^{2}_{\text{res}}=1)$. A three-level diet treatment $T$ was assigned cyclically to animals but purposely left inactive in the data-generation equation, so that the global null H_0:TH\_0:T has no effect is true by construction. We examined both a balanced layout (all 72 animals observed) and an unbalanced layout in which 20% of animal records were deleted at random, reflecting typical field attrition. For each layout, we generated 1,000 replicate datasets and applied two $F$-tests for the diet factor: (i) Exact-$\mathrm{df}$ test—numerator $\mathrm{df}$ =2=2 (levels 1-1) and denominator $\mathrm{df}$ equals the within-animal residual $\mathrm{df}$ obtained from the projector–rank decomposition. (ii) Naive one-way ANOVA—treats all observations as independent, using the pooled residual mean square and denominator $\mathrm{df}=N-3$.
The results are tabulated in Table 6.
Layout
Exact-$\mathrm{df}$
empirical size
Naïve
empirical size
Balanced (72 animals) 2.7% 0.3%
20% animals missing 7.7% 1.0%
Table 6: Empirical size comparison of exact $\mathrm{df}$ and naïve approaches for the nested animal model.
The projector–rank test tracks the nominal 5% level remarkably well—slightly conservative in the balanced case (2.7%) and mildly liberal (7.7%) when one-fifth of records are missing, both within the Monte Carlo margin of error for 1,000 replicates. In stark contrast, the naïve analysis is ultra-conservative, rejecting only 0.3% of the time when balanced and 1% after data loss. Ignoring the nesting hierarchy collapses sire- and dam-level variability into the denominator, inflating the error variance and the associated $\mathrm{df}$; the resulting test lacks power and dramatically understates the type-I error. Juxtaposing this finding with the split-plot experiment reveals a symmetrical danger. Misallocating $\mathrm{df}$ can swing either liberal (false discovery) or conservative (missed discovery) depending on the design: liberal when whole-plot variation is ignored, conservative when higher-level clustering is disregarded. The projector–rank partition theorem navigates both pitfalls automatically by aligning each factorial contrast with the stratum where its random variability actually resides. Moreover, its performance is robust to substantial imbalance—here, 20% missing animals—something traditional Satterthwaite or Kenward–Roger approximations often struggle with. For practitioners in animal science, genetics, and any multi-stage sampling context, adopting the exact $\mathrm{df}$ is therefore crucial to maintaining valid inference and avoiding costly type-I or type-II errors.
As a third experiment, we now look at information-retention and power optimization in blocked fractional factorial designs that showcases benefits of the projector–rank framework. When an experimenter must fractionate a $2^{k}$ factorial and accommodate a blocking constraint (e.g., oven batches or day-to-day shifts), how should generators be chosen? Classical criteria (resolution, aberration) ignore blocking; conversely, blocking handbooks treat only full factorials. The projector–rank machinery supplies quantitative guidance through the $\mathrm{df}$-retention ratio $\rho(\overline{E})$. We demonstrate that maximizing $\rho$ leads to measurably higher power for detecting main effects while using the same number of runs. Table 7 compares a $2^{5-1}$ fractional factorial experiment with traditional and $\rho$-optimized designs.
Feature Design A Design B
Runs 16 ($2^{5-1}$) 16 (same)
Factors AA, $B$, $C$, $D$, $E$ AA, $B$, $C$, $D$, $E$
Generators $I=ABCDE$ $I=ABCD$, ACEACE
Blocking
2 day-batches of 8 runs
(block generator: BCDEBCDE)
same run-budget, block generator
chosen by integer-programming
to maximize $\sum_{E}\rho(\overline{E})$
Main-effect $\rho$ $A,B,C=0.5$ $A,B=1$, $C=0.75$
Alias loss $\delta$ 5 $\mathrm{df}$ lost 2 $\mathrm{df}$ lost
Table 7: Compact summary of traditional (A) vs $\rho$-optimized (B) $2^{5-1}$ designs.
Design B retains two extra $\mathrm{df}$ for the three primary factors $A$, $B$, $C$ at no extra cost. The simulation protocol involves the following model: $y=\mu+\beta_{A}A+\beta_{B}B+\beta_{C}C+\text{Block}+\varepsilon$, with block $\sim\mathcal{N}(0,\sigma^{2}_{b}=2)$ and $\varepsilon\sim\mathcal{N}(0,1)$. The true effects include $\beta_{A}=\beta_{B}=1$, $\beta_{C}=0$ (so power and type-I error can both be assessed). The analysis models involve (i) mixed model using projector–rank exact $\mathrm{df}$ (denominator = 7 for all tests) and (ii) mixed model with Kenward–Roger $\mathrm{df}$ (software default). We employ the following metrics: Power for $A$, $B$ at $\alpha=0.05$, type-I error for $C$, and width of 95% CI for $\beta_{A}$. 10,000 Monte Carlo replicates were run for each design. The results are tabulated in Table 8.
Design & method Power (A) Power (B) Type-I (C) Mean CI-width (A)
A–exact $\mathrm{df}$ 0.42 0.41 0.053 1.87
A–KR approx 0.36 0.35 0.041 2.05
B–exact $\mathrm{df}$ 0.61 0.60 0.051 1.47
B–KR approx 0.55 0.54 0.040 1.63
Table 8: Comparison of statistical properties under exact $\mathrm{df}$ and Kenward–Roger (KR) approximation for Designs A and B.
It is seen that simply swapping generators to maximize $\rho$ recovers 3 lost $\mathrm{df}$, translating to $\approx$ 20% narrower confidence intervals and $\approx$ 60% power instead of $\approx$ 40% for main-effect detection. Even when Kenward–Roger is applied, Design B still outperforms Design A—but exact $\mathrm{df}$ always deliver the best power/coverage because they avoid the slight conservatism Kenward–Roger introduces. The partition-theorem machinery is not only about testing with the right denominator; it also guides design construction. By quantifying which contrasts are sacrificed by fractionation + blocking ($\delta$), the experimenter can choose generators that maximize information where it is most needed, without adding runs.
Finally, we conduct experiments to assess the computational-efficiency of the proposed method. To quantify how much faster the projector–rank calculus delivers denominator $\mathrm{df}$ than a popular simulation-based alternative (for example, parametric bootstrap) when sample size $N$ grows to the scale encountered in modern high=throughput experiments. The following computing environments and general settings were employed. The hardware comprised AMD Ryzen 7 5700U CPU 1.8 GHz; 16 GB RAM; no GPU acceleration. We employ Python 3.11 and standard libraries such as NumPy, SciPy; benchmarked with time.perf_counter(). Each experiment was run 30 times; reported figures are medians to mute OS-level jitter ($3\%$ SD across runs). We let $N=1,000;10,000;100,000$, representing small, moderate, and large data sets. Ten equal-sized blocks per $N$ (to mimic lab batches) were used. Block membership was encoded as integer labels $0,\dots,9$.
We compared two methods for estimating exact $\mathrm{df}$. The projector–rank method builds the block-mean projector $P=\operatorname{diag}(1/n_{j}),Z,Z^{\prime}$, applies it to a random contrast vector, and computes the resulting one-column Gram rank. This approach is computationally efficient, with $O(N)$ complexity for forming block means, and requires only $O(N)$ memory through the use of streamed averages. It is implemented in pure NumPy using two core steps: (i) block sums are computed via np.bincount, and (ii) the projected means are assigned using means[labels]. Importantly, no $N\times N$ matrix is ever formed. In contrast, the parametric bootstrap approach generates $B=100$ random $F$-ratios under the null and matches their moments to approximate the denominator degrees of freedom—a standard shortcut for small-sample mixed models. This method has higher complexity, $O(BN)$, since each resample requires fresh memory allocation. It relies on SciPy’s stats.f.ppf along with NumPy random draws. The choice of $B=100$ follows the minimal resampling threshold often recommended in software documentation. The bootstrap baseline is deliberately favorable: Kenward–Roger can be 2–4 $\times$ slower than 100 bootstrap draws because of repeated matrix inversions; thus the reported speed differences are lower bounds.
The following procedure is employed as the timing protocol. For each $N$:
  1. 1.
    Input generation: Draw a single random double-precision vector $v\sim\mathcal{N}(0,1)$ of length $N$; generate the block-label vector of equal size. Input generation is excluded from timing to isolate algorithmic cost.
  2. 2.
    Timing window: Start perf_counter; run either the projector routine (compute means, project $v$, rank) or the bootstrap loop (100 simulations and $F$-ratio computation); stop the clock.
  3. 3.
    Memory: Peak RAM for the projector version never exceeded $0.9\times N$ doubles ($\approx$ 0.75 MB per 100, 000 rows) because only two length-$N$ arrays exist concurrently. Bootstrap briefly holds three length-$N$ arrays per replicate, peaking near $B\times N$ doubles (64 MB at $N=100,000$).
The results are tabulated in Table 9 (median of 30 runs).
NN Projector–rank time (s) Bootstrap time (s) Speed-up factor
[…]
Table 9: Execution time and speed-up factor for projector–rank method versus bootstrap as sample size NN increases.
Times grow linearly with $N$ for the projector method (a ten-fold increase from $N=10^{4}$ to $10^{5}$ raises CPU time only 6 ms, matching the cost of one extra bincount). Bootstrap costs scale almost exactly with $B\times N$ and dominate beyond $N\approx 10^{4}$. The exact $\mathrm{df}$ calculation is effectively an $O(N)$ streaming average; wall-clock time remains sub-10 ms even at $N=10^{5}$. Bootstrap times grow from hundredths to more than half a second—and would exceed 6s at $N=10^{6}$ if $B=100$ were kept. Projector code maintains two vectors regardless of $B$; bootstrap scales as $O(BN)$, becoming memory-intensive in extreme high-throughput settings (e.g., genomic scans, sensor data). The projector approach yields deterministic integer $\mathrm{df}$; bootstrap introduces Monte Carlo error that must be re-estimated if $B$ changes.
While split-plot simulation illustrated liberal versus exact size, nested animal simulation showed the opposite risk, i.e., over-conservatism. The fractional-factorial study demonstrates a design-selection application, confirming that higher $\rho$ translates to tangible gains in estimator precision and test power. Lastly, for modern data volumes the projector–rank framework is not only more statistically valid than approximation-based $\mathrm{df}$, but also orders of magnitude faster and lighter on memory, making it attractive for pipeline automation and large-scale screening where thousands of mixed-model fits must run overnight.
6 Concluding remarks
In this section, we discuss some computational aspects applicable in practical settings and provide concluding remarks. A stratum is any collection of experimental units that shares the same source of random variation (variance component) and is averaged over in the randomization process before lower-level treatments are applied. Typical sources are tabulated in Table 10.
Random component Typical label Examples
Single random factor Block, Batch, Litter
“Block” in RCBD; “Batch” in drug assay
Interaction of
random factors
Block $\times$ Day Two-way randomized complete block
Nested factor Dam Sire $\rightarrow$ Dam $\rightarrow$ Animal hierarchy
Cross-classified unit
School &
Classroom
Multi-site trials with teachers nested
in schools but students crossed
Table 10: Types of random components in experimental designs.
If a factor is fixed, it does not generate its own $\mathbf{P}_{s}$; it merely defines contrasts inside the lowest stratum that still carries variability. Interactions that contain at least one random component inherit the higher of the two strata. For a stratum $s$ with $q_{s}$ distinct unit labels, let $\mathbf{Z}_{s}$ be the $N\times q_{s}$ incidence matrix: $(\mathbf{Z}_{s})_{ij}=1$ if observation $i$ belongs to unit $j$; $n_{j}$ the replication within unit $j$. The idempotent projector $\mathbf{P}_{s}=\mathbf{Z}_{s}\,\left(\mathbf{Z}_{s}^{\prime}\mathbf{Z}_{s}\right)^{+}\mathbf{Z}_{s}^{\prime}=\operatorname{diag}\left(\tfrac{1}{n_{u(i)}}\right)_{i=1}^{N}\mathbf{Z}_{s}\mathbf{Z}_{s}^{\prime}$ replaces each response by the mean of its stratum-$s$ unit. In balanced layouts $\mathbf{Z}_{s}^{\prime}\mathbf{Z}_{s}=n_{s}\mathbf{I}$ and $\mathbf{P}_{s}=\frac{1}{n_{s}}\mathbf{Z}_{s}\mathbf{Z}_{s}^{\prime}$. For sparse or very large $N$, compute an integer vector unit[i] listing the stratum-$s$ label for observation $i$. Compute means mu = np.bincount(unit, y)/np.bincount(unit) and projected values yhat = mu[unit]. To apply $\mathbf{P}_{s}$ to any numeric vector $\mathbf{v}$, replace ‘y’ in the snippet by $v$. Sparse-matrix libraries (e.g., scipy.sparse) can build $\mathbf{Z}_{s}$ and perform $\mathbf{P}_{s}X$ in $O(\text{nnz}(\mathbf{Z}_{s}))$ when multiple columns are needed. For nested random factors, we build one projector per level of nesting, e.g., in Sire $\rightarrow$ Dam $\rightarrow$ Animal we get $\mathbf{P}_{\text{Sire}}$ and $\mathbf{P}_{{\text{Dam}|\text{Sire}}}$. The residual (Animal) projector is implicitly $\mathbf{I}-\sum\mathbf{P}_{s}$. For crossed random factors, e.g., if Block and Day are both random and fully crossed, each generates its own projector; their interaction Block $\times$ Day is a third random component with $\mathbf{P}_{\text{Block}\times\text{Day}}=\tfrac{1}{n_{bd}}\mathbf{Z}_{\text{Block}\times\text{Day}}\mathbf{Z}^{\prime}_{\text{Block}\times\text{Day}}$. The rule of thumb is “One variance component $\Rightarrow$ one projector.”
Once $\mathbf{P}_{s}$ is ready, projectors are constructed. In sparse, unequal-replication cases this is cheaper via averaging than by full $N\times N$ matrices. Factorial contrasts are enumerated using tensor products of orthogonal polynomial contrasts or Helmert contrasts and account for aliasing if fractionated. Then, ranks are computed, where we multiply small $k\times k$ Gram matrices $\mathbf{X}_{\overline{E}}^{\!\prime}\mathbf{P}_{s}\mathbf{X}_{\overline{E}}$ (dimension $\leq$ $\mathrm{df}$ of the effect) and take ordinary rank. Complexity of this step is $O\left(\sum_{E,s}\operatorname{df}(E)^{3}\right)$, trivial relative to fitting the mixed model. As a sanity check, we can verify $N-1=\sum_{E,s}\text{$\mathrm{df}$}_{E,s}$. Any shortfall points to data omissions or coding errors.
By proving the $\mathrm{df}$ partition theorem we showed that every mixed, blocked, fractionated, or unbalanced design possesses an exact, integer $\mathrm{df}$ decomposition achievable via rank calculations. The theorem’s simultaneous action of stratum projectors and level-permutation symmetries replaces an existing case-by-case derivations with a single algebraic identity. Exact numerator and denominator $\mathrm{df}$ restore nominal error rates in $F$-tests where Satterthwaite and Kenward–Roger approximations often mislead. The new retention, deficiency, and inflation indices ($\rho,\delta,\alpha$) quantify how blocking, missing plots, or fractional generators erode information-something that existing resolution metrics cannot do in multi-stratum settings. Quantifying the lost $\mathrm{df}$ lets us decide whether extra runs are worth the cost. REML score equations simplify when the $\mathrm{df}$ of each stratum are known exactly; closed-form starting values accelerate convergence. Exact $\mathrm{df}$ tables demystify why, for example, the $AB$ interaction in a split-plot draws its denominator from sub-plot error instead of whole-plot error.
From a computational standpoint, $\mathrm{df}$ depends only on small Gram-matrix ranks, making the results directly compatible with statistical packages (R/Python). Adding them to existing lmer or mixedlm wrappers would offer analytic $\mathrm{df}$ alternatives to numeric approximations. Replacing $\mathbf{P}_{s}$ with working-weight projectors extends the rank identity to non-Gaussian outcomes, enabling analytic $\mathrm{df}$ for logistic split-plots or Poisson strip-plots, both currently Monte Carlo-based. These directions may yield new insights.